### v3

# AURA: Autonomous Upskilling with Retrieval-Augmented Agents

Abstract
Designing reinforcement learning curricula for agile robots traditionally requires extensive manual tuning of reward functions, environment randomizations, and training configurations. We introduce AURA (Autonomous Upskilling with Retrieval-Augmented Agents), a schema-validated curriculum reinforcement learning (RL) framework that leverages Large Language Models (LLMs) as autonomous designers of multi-stage curricula. AURA transforms user prompts into YAML workflows that encode full reward functions, domain randomization strategies, and training configurations. All files are statically validated before any GPU time is used, ensuring efficient and reliable execution. A retrieval-augmented feedback loop allows specialized LLM agents to design, execute, and refine curriculum stages based on prior training results stored in a vector database, enabling continual improvement over time. Quantitative experiments show that AURA consistently outperforms LLM-guided baselines in generation success rate, humanoid locomotion, and manipulation tasks. Ablation studies highlight the importance of schema validation and retrieval for curriculum quality. AURA successfully trains end-to-end policies directly from user prompts and deploys them zero-shot on a custom humanoid robot in multiple environments —capabilities that did not exist previously with manually designed controllers. By abstracting the complexity of curriculum design, AURA enables scalable and adaptive policy learning pipelines that would be complex to construct by hand. aura-research.org
I Introduction
Curriculum reinforcement learning (RL) [narvekar2020curriculum, Li2024RLbipedal, anymalparkourlearningagile, fourierlatentdynamics] enables robots to master complex skills by decomposing tasks into progressively harder subtasks [rl_curriculum]. This approach has proven to be effective in domains like agile locomotion [sciencesoccer, real_world_digit, massive_rl, learningtorquecontrolquadrupedal], where single-stage learning can struggle with sparse rewards or high-dimensional exploration spaces. By splitting the learning process, agents are guided through increasingly challenging environments, improving sample efficiency and convergence.
However, designing effective multi-stage curricula remains a significant bottleneck [jang2018combining, nguyen2021robots]. Multi-stage curricula require coordinated changes across reward shaping, randomization, simulation fidelity, and optimization; as the number of stages increase, designing transitions, verifying stability, and tuning parameters scales combinatorially (the “curse of dimensionality” [li2025causally]) and becomes brittle to human error and heuristic bias [sayar2024diffusionCurriculum, soviany2022curriculum]. Moreover, failures at any stage often derail the entire training process, making automation critical for scaling curriculum RL.
Large language models (LLMs) offer a compelling alternative to optimization-based approaches [hu2020learning]. Their high-level reasoning and generalization capabilities have demonstrated success in robotic planning, perception, and code generation [brohan2022rt, ahn2022can, liang2023code, sun2024prompt, mower2024ros, mon2025embodied]. While LLMs are not directly deployable on real-time control loops due to latency constraints [team2025gemini], LLM’s have shown potential in designing training schemes for real-time robot control policies [ma2023eureka, ryu2024curricullm, ma2024dreureka]. However, current LLM-based RL pipelines inefficiently use computation on launching parallel environments to deal with malformed generations and fail to learn from experience across tasks.
What is missing is a principled framework for transforming high-level prompts into reliable, executable RL training pipelines that improve with experience. Such a system should (1) ensure syntactic and semantic validity before execution, (2) use past experience to optimize future performance, and (3) modularize the pipeline into verifiable components [Erak_2024]. Without these capabilities, LLMs are limited to more fragile, trial-and-error reward design rather than scalable curriculum generation.
AURA (Autonomous Upskilling with Retrieval-Augmented Agents) addresses these bottlenecks with three key ideas. (i) A typed YAML schema captures curricula, reward functions, randomization, and training hyperparameters; structured LLM outputs are statically validated before a single GPU cycle is spent. (ii) A team of specialized LLM agents works in collaboration to enable modular, schema-compliant curriculum generation. (iii) A Retrieval-augmented generation (RAG) module with a vector database (VDB) stores prior task specifications, curricula, and rollout evaluations. This database supports experience-informed generation by enabling agents to condition on successful prior workflows, improving both accuracy and generalization.
We evaluate AURA on a suite of humanoid locomotion and deploy the resulting policies zero-shot on a custom, kid-sized humanoid. Compared to curriculum RL baselines and recent LLM-crafted reward pipelines, AURA is capable of generating context-rich, high-dimensional curricula that produce successful zero-shot policies deployable on real humanoid hardware. Our contributions are:
1.
AURA: A fully agentic, retrieval-augmented framework that turns a natural-language prompt into a hardware-deployable controller policy.
2.
Curriculum compiler, schema-validation: A typed YAML schema that provides an LLM-friendly interface for defining reward terms, domain randomizations, and training configurations. Static validation of the YAMLs enables descriptive error messages for generation retries, without wasting compute on failed environment launches.
3.
Experience-aware, self-improving, curriculum generation agent: Specialized LLM agents query a vector database of past runs and evaluations, select relevant training files, generate high-level curricula, and refine them into executable multi-stage specifications. This feedback-driven loop promotes curriculum quality and training stability over many iterations.
4.
Zero-shot deployment: Experiments showing AURA’s policies transfer zero-shot to a custom kid-sized humanoid, Berkeley Humanoid, Booster T1, [berkeleyhumanoid], and manipulators, showing end-to-end prompt-to-policy ability.
II Related Work
II-A Large Language Models in Robotics
Early efforts to link natural language to robot control and task-planning frameworks translate natural-language commands into symbolic action graphs, as demonstrated by SayCan and Code-as-Policies [ahn2022can, liang2023code]. Recent Vision–Language–Action paradigms, such as RT-1/2, , Octo, and OTTER, [brohan2022rt, zitkovich2023rt, pi0, octo, otter], demonstrate that pretraining on large-scale robot data yields policies that generalize across many robotic skills. Further work has extended this to cross-domain generalization, deformable object manipulation, and long-horizon household tasks, exemplified by Gemini Robotics and embodied LLMs [mon2025embodied, team2025gemini]. Other lines of work have explored LLM-driven planning and language-augmented reasoning for robotic control pipelines [sun2024prompt, mower2024ros, mon2025embodied, llm_robotics_survey, ai_review_humaniod, llm_robot_task_plan, rho2024language]. Despite their semantic flexibility, these methods continue to rely on hand-tuned, low-level controllers and standard actuation assumptions, with LLMs typically external to the closed control loop [llm_robotics_survey, ai_review_humaniod]. RoboGen addresses these limitations through decomposing long-horizon tasks into skills, which are tuned or trained by LLMs [auerbach2014robogen].
II-B Curriculum Reinforcement Learning
Reinforcement learning often struggles with sparse rewards and long‐horizon problems, where meaningful feedback is rare or delayed over many timesteps [discover]. Curriculum RL addresses sparse-reward and long-horizon problems by exposing agents to a progression of gradually more difficult tasks [narvekar2020curriculum, rl_curriculum, soviany2022curriculum]. Surveys and benchmarking studies note that practical curriculum design remains heuristic and sensitive to context-specific particulars [soviany2022curriculum]. Reverse-curriculum approaches [reverse_curriculum_rl, tao2024reverse] employ a staged training process that progressively increases the task’s difficulty from previous successful demonstrations, thereby enhancing exploration and sample efficiency. Domain-randomization and curriculum-based RL have enabled real-world legged locomotion and manipulation in challenging environments [Li2024RLbipedal, anymalparkourlearningagile, fourierlatentdynamics, sciencesoccer, real_world_digit, massive_rl, learningtorquecontrolquadrupedal, zhuang2024humanoidparkourlearning, quadsoccershoots, dribblebot, nimbro_2022, adu2023exploring].
II-C LLM-guided RL and Curricula
Emerging systems such as CurricuLLM [ryu2024curricullm] and Eureka [ma2023eureka, liang2024eurekaverse] utilize LLMs to generate reward functions for RL training. Eureka uses evolutionary search and prompts LLMs in parallel to directly generate reward code for various tasks in simulation. DrEureka extends Eureka to generate domain randomizations and consider hardware safety, demonstrating real-world locomotion tasks on a Unitree Go1 quadruped robot [ma2024dreureka]. CurricuLLM [ryu2024curricullm] similarly uses evolutionary search and directly generates code, but focuses on curriculum generation and humanoid locomotion, showing real-world success on the Berkeley Humanoid [berkeleyhumanoid].
These frameworks require sampling a large number of training runs to obtain a viable or high-performing policy, relying on LLM stochasticity and brute-force search. RAG and VDBs are approaches to enhance the LLM knowledge beyond what the LLM has been trained on [liu2025hm], which AURA utilizes to construct effective curricula.
Positioning of this work: AURA unifies curriculum generation, domain randomization, and training configuration within a schema-validated, retrieval-augmented loop that learns from prior runs. This yields consistent, deployable specifications with higher first-attempt launch rates and removes the need for post hoc sampling or manual curation.
III Methods
III-A Problem Setup and Curriculum Formalism
We define a curriculum as an ordered sequence of training stages, , where each stage is comprised of a reward , domain-randomization distributions over environment parameters , a training configuration , and a promotion criterion that governs advancement. The curriculum stages are modeled as an MDP , [sutton1998reinforcement, puterman2014markov] with denoting randomized physics properties [tobin2017domainrandomization, muratore2021data]. AURA generates for each stage as schema-compliant YAML files that are statically validated and compiled into an MJX-based RL pipeline [todorov2012mujoco, brax2021github]. We use this notation to describe generation, validation, training, and iteration.
III-B Inputs to Curriculum Generation
To instantiate a curriculum from a prompt, AURA consumes four inputs: (i) a natural language task description specifying the desired behavior; (ii) a structured robot specification: joint names, semantic labels (e.g., “left foot contact”), available sensors, and actuator limits; (iii) an MJCF simulation model encoding robot kinematics, dynamics, and contact geometry; and (iv) a Python MJX environment defining observations, actions, physics parameters, and episode termination conditions.
III-C Retrieval-Augmented High-Level Planning
AURA begins by leveraging LLMs and a vector database of prior curricula and outcomes to generate a high-level plan. The user-provided task description and robot specification are first passed into a query LLM, which synthesizes them into a structured retrieval query. This query is then encoded with OpenAI’s text-embedding-3-large [openai_text_embedding_v3] and used to perform cosine-similarity search over the VDB (Pinecone). From the top- most relevant prior curricula, the database returns reward, randomization, and training files, along with user feedback. A selector LLM then chooses the single most relevant example from this set, based on both user feedback and curricula relevance to the current task; this example is included in the high-level planner’s context for generating the new curricula.
Conditioned on this selected example and the user task specification, a high-level planner decides the number of stages and produces a natural-language plan for each one. The plan describes ideas for reward components, domain randomizations, and stage-specific training hyperparameters (e.g., terrain type, disturbance level, PPO hyperparameters, etc.).
The VDB enables two modes of AURA. (1) AURA Blind where the VDB is initialized as empty for the first iteration and AURA must design a reward from scratch, and (2) AURA Tune where the VDB is initialized with human-expert rewards from a related task and must tune that reward for the current domain.
III-D Stage-Level Specification and YAML Generation
Each stage plan from the high-level planner is given to an independent stage-level generator which translates the stage plan into structured YAML files. The stage-level generator takes as input: (i) the user task description; (ii) the natural-language stage description produced by the high-level planner; (iii) the available reward function library; (iv) state variables and sensor signals exposed by the robot environment; (v) the selected prior curriculum example retrieved from the VDB (chosen from the top- retrieval set); (vi) context from previously generated stages in the current workflow; and (vii) the reward/schema format specifications.
Each stage description is translated into three human-readable, schema-compliant YAML files: Reward YAML defines the differentiable terms, coefficients, and aggregation logic with explicit sensor/state dependencies; Randomization YAML defines the target parameters, distributions, and activation conditions; and Training Config YAML defines the PPO hyperparameters, episode budgets, checkpointing cadence, normalizations, etc.. These files provide the abstract specifications needed for execution without requiring direct generation of JAX code.
III-E Schema Validation and Compilation
Directly emitting JAX/MJX code from LLMs can be brittle in practice. AURA compiles only after static validation against typed schemas governing the generated workflow, reward, randomization, and training files. Validation enforces: (i) type conformance (int, float, bool, vector, etc.); (ii) structural schema compliance; (iii) reference integrity (all state variables and sensors referenced in exist in the MJX environment); and (iv) mathematical well-formedness of reward expressions. Expressions are built from a function registry (e.g., fn.NORM_L2 producing JAX arrays) and must compile symbolically under MJX. A generation retry is triggered whenever schema validation detects an error. The schema validation generates a descriptive error message that includes the type of error, content of what caused the error, and sometimes a recommendation for how the error can be fixed. This error message, along with the previously generated (but invalid) YAML files, are fed back into the next attempt to guide corrections and avoid repeating mistakes; we set a maximum of five retries per stage. Once validated, YAMLs are compiled into MJX code.
III-F Staged RL Training Loop
Compiled stages are trained with PPO across parallel simulation environments, each running for the episode limit specified in . Policies from one stage initialize the next, enabling progressive skill learning.
III-G Feedback and Iteration
Following each stage, an automated feedback module analyzes rollouts and training signals under a task-specific rubric (e.g., success rate, stability and energy, reward-component attribution). Its recommendations may adjust reward terms, randomizations, or hyperparameters for subsequent stages within the same iteration if the curriculum training has not yet completed all stages.
An iteration of our framework is defined as the complete execution of: (i) VDB-backed retrieval and high-level planning; (ii) YAML generation and static validation for all stages; (iii) PPO training across stages; with (iv) automated feedback analysis between each stage. After the final trained policy is deployed in simulation or on hardware, the user will evaluate the policy either based on the simulation video rollout or hardware performance, then provide user feedback: 2–4 sentences in natural language judging the policy’s behavior and quality based on the video rollout (eg. ”the policy survives well, but could track linear velocity better”). User feedback is attached to the files of the just-trained curriculum and inserted into the VDB alongside the YAMLs, metrics, and rollouts. Any training that incorporates user feedback begins a new iteration , which again starts from VDB retrieval conditioned on the now-augmented database. Human input therefore does not mutate an ongoing run; it shapes the next planning phase via retrieval.
III-H Modular LLM Collaboration
Hallucination mitigation: AURA limits LLM hallucinations through three mechanisms: (i) retrieval grounding—the planner is context-seeded by a single example selected from the top- VDB results; (ii) schema constraints—the model emits only typed YAML using a restricted operator registry (e.g., fn.NORM_L2) rather than arbitrary code; and (iii) consistency checks—static validation enforces type/structure and reference integrity to environment signals, followed by MJX compile checks and a capped regenerate-on-error loop (max retries). Together, these mechanisms reject unsupported symbols and incoherent references, improve first-try success, and reduce off-distribution generations.
IV Experiments
IV-A Simulation Environments
We present simulation experiments on a variety of tasks on three humanoid robots and a robot arm manipulator, all example AURA trained policy rollouts can be seen in Fig. 1. All AURA experiments are trained in MuJoCo-MJX and are evaluated over 1024 parallel environments on 5 seeds. CurricuLLM simulation results are reported based on their results in Isaac Lab.
IV-A1 Custom Humanoid
The custom humanoid locomotion environment exposes the velocity command, joint positions, projected gravity vector, last action, phase, and max foot height as observations and actuator gains, friction coefficients, body mass, COM position, foot contact geometry, initialization configuration, and roughness as domain randomization parameters. The simulation setup was adapted from the model in [bruce_parallel_sim]. AURA is run for five iterations, with each iteration generating a new curriculum. Custom humanoid locomotion is evaluated with AURA Tune, which starts with the Berkeley Humanoid reward and adapts it into a curriculum that works on the custom humanoid hardware, and AURA Blind, which generates a curriculum from scratch. All environments are run at Hz, and the curriculum is limited to M training timesteps.
| Episode Survival Length | Linear Vel. Tracking | Pushing Cube Success % | |
| AURA | |||
| Playground | ** | ||
| CurricuLLM* |
IV-A2 Berkeley Humanoid
For the Berkeley humanoid task, AURA uses the MuJoCo Playground’s Berkeley Humanoid environment. Berkeley Humanoid’s MuJoCo Playground policy is trained as detailed in [mujoco_playground_2025]. The environment is run at 50Hz and both policies train over M steps.
IV-A3 BoosterT1 Humanoid
For BoosterT1 locomotion, AURA is trained and evaluated in the MuJoCo Playground BoosterT1 simulation environment, is run at 50Hz, and trains for M steps. BoosterT1’s MuJoCo Playground policy is trained as detailed in [mujoco_playground_2025].
IV-A4 UR5e Manipulator
The environment observations and domain randomizations are consistent with MuJoCo Playground’s manipulator environment for the Franka Emika Panda (but with a UR5e robot). The environment frequency is Hz, and the curriculum is limited to M training steps. This environment is used for two tasks. In the pushing cube task the robot must push a block to a target position both of which are randomized in a cm by cm region (following CurricuLLM experiments). To be considered a success the center of the cube must be within cm of the target point. The stacking cube task begins with a cube in the gripper and the robot must place the cube on top of another block which is randomized in a cm by cm region.
| AURA | AURA w/o schema | AURA w/o VDB | AURA single-agent | CurricuLLM∗ | Eureka∗∗ | |||
99 % | 47 % | 38 % | 7 % | 31 % | 12 / 49 % |
IV-B Experimental Setup
We evaluate two aspects of AURA and baselines:
(1) Generation Success Rate. In Table II, we report the success rates of AURA, its ablations, CurricuLLM, and Eureka. For AURA, we measure curriculum-level success. A run is considered successful only if all stages (rewards, domain randomizations, and training configs) in the generated curriculum successfully launch GPU-accelerated training. AURA performs a single launch per task without parallel sampling. We also compare AURA with its ablated variants: AURA without a VDB, AURA without schema validation (no retires), and AURA using a single LLM for both high-level planning and YAML generation. Eureka and CurricuLLM launch multiple environments in parallel. We report the fraction of successful environment launches out of all generated rewards. Unlike AURA, success is measured one stage at a time rather than requiring a full curriculum to succeed.
(2) Training and Deployment Benchmarking. To evaluate policy quality, we run five iterations of curriculum generation and training using AURA, AURA Blind, AURA Tune, CurricuLLM, and a human-designed baseline. Each trained policy is evaluated with random seeds each across randomized environments. Evaluation reward functions, domain perturbations, and task variations are hidden from all methods to simulate realistic zero-shot deployment. For locomotion, survival episode length is the average number of steps survived without the humanoid falling (max 3000). Linear velocity tracking score is calculated per episode as:
Where and are the commanded and actual robot velocities at timestamp and is the number of timsteamps the policy survives for.
Selected final policies are transferred to hardware for qualitative testing and demonstration.
IV-C Simulation Experiment Evaluations
IV-C1 Custom Humanoid
Fig. 3 summarizes evaluation during simulation deployment across iterations, showing average survival and linear velocity tracking scores. Both AURA variants improve noticeably over the human baseline on survival and linear-velocity tracking, which shows that AURA can generate an effective curriculum from scratch and can adapt existing rewards into effective curriculums. AURA Tune and AURA Blind both iteratively improve over five iterations, reasoning about reward signals and user feedback. AURA Blind’s curriculum did not generate or use any rewards related to foot swing or foot phase tracking, leading to a more inconsistent and asymmetrical gait, but better simulation performance compared to AURA Tune. The locomotion task prompt does not specify stylistic information about the gait, so this is a reasonable result. AURA Tune, which is initialized with MuJoCo Playground’s rewards, kept existing reward terms for phase tracking, maintaining a consistent gait while still improving over the base reward.
IV-C2 Berkeley Humanoid
For the Berkeley Humanoid results shown in Table I, AURA’s episode survival length is similar to the MuJoCo Playground baseline, showing that both survive nearly the full 3000-step horizon. The key difference is in linear velocity tracking, where AURA achieves a significantly better score, reflecting more precise command following. While Playground training schedules M steps of flat terrain then M steps on rough terrain, AURA split training into M on flat terrain with mild perturbations, M on rough terrain with mild perturbations, and M on rough terrain with heavier, more frequent pushes, while progressively expanding the command range up to the 1.0 m/s and 1.0 rad/s caps. AURA also made the velocity following reward more strict and increased reward emphasis on feet-phase and action-rate, resulting in better velocity command following and smoother, more consistent stepping.
IV-C3 BoosterT1 Humanoid
Training with the MuJoCo Playground’s BoosterT1 reward directly resulted in a policy with jerky motion and small, high-frequency shakes during locomotion. AURA was told this as feedback, which led it to generate new reward terms which penalize the squared differences across consecutive time steps for actions and for velocities, resulting in a visibly smoother policy. Heavier domain randomization also helped improve performance, raising episode length from to steps and linear-velocity tracking score from to .
IV-C4 UR5e Manipulator
For the UR5e cube pushing task, AURA begins with MuJoCo Playground’s Franka Emika Panda environment reward which is not sufficient to train a policy in the UR5e environment as shown in Table I. AURA adapted the reward into a curriculum where the first stage has no positional domain randomization and reshaped the gripper–to-cube reward term to provide a denser reward signal over larger distances, encouraging progress toward the cube rather than only near contact. In the second stage, AURA reintroduced the positional domain randomization and increased the weight of the box to target reward. This led to a final policy which could consistently push the cube to the target location with over a % success rate.
For the cube stacking task AURA adapted the push task curriculum, dramatically decreasing the gripper-to-cube reward because the cube starts in the gripper, and emphasizing the cube-to-cube reward term. AURA achieved a mean task success rate of 72.7%.
IV-D Real-World Experiments
Hardware Setup: We use a custom m-tall agile humanoid with 5-DoF legs and 3-DoF arms [bruce]. RL policies run at Hz with sensors and actuators at Hz. The end-to-end policy uses no estimators or low-level control beyond servo internal impedance controllers.
Fig. 1 shows a policy generated by AURA deployed on real hardware in a zero-shot setting. The outdoor locomotion policy shown in the figure is trained by AURA Tune, while the robot traversing a step is trained by AURA Blind. In response to a user prompt requesting outdoor-capable locomotion, AURA automatically generated a four-stage curriculum progressing from flat terrain with low randomization to rough terrain with high randomization, diverse velocity commands, and significant external perturbations. The resulting policy demonstrated robust zero-shot generalization when deployed outdoors, successfully handling uneven terrain and maintaining consistent gait tracking. The robot achieved walking at and average gait frequency of . It withstood substantial perturbations, recovering from lateral pushes up to and from being physically pushed off a high platform without falling, which surpassed the performance of existing manually tuned controllers [bruce].
IV-E Generation Success and Efficiency
IV-E1 Curriculum Generation
Table II presents the success rates of training runs for AURA, its ablated variants, and baseline methods including CurricuLLM and Eureka. AURA achieves a success rate, outperforming all other methods and demonstrating the benefit of schema validation and RAG which allow for consistent generation of curriculums with hundreds of parameters and sometimes over ten unique reward terms.
IV-E2 Computation Efficiency
AURA requires only a single training launch per stage, since schema validation and VDB-guided retrieval ensure that nearly all generated curricula are executable on the first training launch. In contrast, Eureka and CurricuLLM each generate multiple training launches per stage—16 and 5, respectively—and retain only the best-performing successful run. As a result, AURA initiates far fewer training runs than either CurricuLLM or Eureka, leading to substantially lower computational cost for multi-stage workflows while still achieving competitive performance.
IV-E3 Multi-Agent LLM Setup
AURA’s use of a multi-agent LLM setup—including a high-level curriculum generation agent and dedicated per-stage agents for reward, domain randomization, and configuration—was critical for producing complete and valid curricula. As shown in Table II, attempting to generate all stages in a single step is highly inconsistent. By decomposing the task into high-level planning and individual stage generation AURA forms smaller, more achievable tasks with higher success rates. The delegation of responsibilities also allows for individual stages to be retried in event of generation failure, rather than all stages.
V Limitations
While AURA automates reward design, domain randomization, and curriculum-driven policy training, a nontrivial amount of human effort is still required. Setting up the MuJoCo simulation environment, defining observation spaces, action spaces, and environment state variables, remains a manual process that can impact policy success. Similarly, deploying the final policy to real hardware requires careful integration and validation.
Human Evaluation: Although AURA’s training pipeline generates curricula autonomously, it still relies on human feedback at the end of each iteration to provide qualitative assessment of the deployed policy. While naively applying Vision-Language Models often fails to accurately describe robot data [robo2vlm, robovqa], recent frameworks demonstrate promising methods for automatically analyzing robot trajectories and delivering feedback [autoeval, rewind]. These could be used in future work to fully automate feedback between iterations.
Exploration and VDB Content Bias: While AURA leverages its VDB to improve training reliability, this dependence introduces potential bias. If the database lacks diversity or contains inaccurate human feedback, subsequent retrieval-augmented generations may reinforce suboptimal patterns. AURA’s curriculum search process has the potential to converge prematurely on a narrow set of solutions. Future work could incorporate explicit exploratory sampling of curricula designs or adversarial selection of prior examples to counteract this VDB content bias.
VI Conclusion
This paper introduces AURA (Autonomous Upskilling with Retrieval-Augmented Agents), a schema-centric curriculum RL framework leveraging multi-agent LLM architectures and RAG. By abstracting complex GPU-accelerated training pipelines into schema-validated YAML workflows, AURA reliably automates the design and refinement of multi-stage curricula, enabling policy training with iterative improvement. Empirical evaluations demonstrate that AURA outperforms baseline methods, achieving a higher generation success rate and superior policy performance across locomotion and manipulation tasks, and effective zero-shot hardware deployment. Real-world validation on a custom humanoid robot highlights AURA’s capability to autonomously produce highly robust policies directly from user-defined prompts. The resulting controller demonstrated stable outdoor locomotion, gait tracking, and disturbance rejection, including recovery from lateral perturbations and vertical drops. Overall, AURA advances the development of scalable, generalizable prompt-to-policy frameworks, demonstrating the ability to substantially reduce the manual effort required in reward design while generating complex, high-performing curricula.
-A Implementation and Testing Details
-A1 Training via Proximal Policy Optimization (PPO)
We train the control policy with PPO, which constrains the policy update to prevent destructive parameter jumps. Let
denote the probability ratio between the new and old policies, and let be the generalized advantage estimate at timestep . The clipped surrogate objective is
where is the clipping parameter (we use unless otherwise noted). During training, we maximize with Adam, using entropy regularization for exploration and a value–function loss with coefficient .
-A2 Score Calculation
During one evaluation run, we launch parallel environments and roll each for a horizon of simulation steps (or until failure). All three metrics are normalized to for direct comparability.
Survival Score.
Let denote the number of steps survived by environment . The Survival Score is the mean fractional episode length
Linear‑Velocity Tracking Score.
For step in environment let be the commanded planar velocity and the robot’s actual planar COM velocity expressed in the local frame. Define the squared tracking error
The YAML exponential_decay entry with corresponds to the per‑step reward
Aggregating over time and averaging across the batch yields the normalized Linear‑Velocity Tracking Score
Summary.
The triplet captures robustness (survival), command‑following fidelity, and gait coordination, respectively, and forms the basis of all quantitative comparisons in the main paper.
-A3 AURA Task Prompts for Training Launch
These five user prompts were used as task inputs to AURA for evaluating its curriculum training-launch success rate across the custom humanoid and Berkeley Humanoid embodiments. For baseline comparisons, we used the prompts provided in each baseline’s open-source repository to generate their respective training files.
-A4 Training-Launch Comparison Details
In the following section, we provide a more detailed description of the environments used to evaluate the CurricuLLM and Eureka training-launch success rates. All environments were evaluated over training launches.
CurricuLLM
CurricuLLM’s training-launch success rate was evaluated on their Berkeley Humanoid environment from [ryu2024curricullm], with all environment scripts and tasks taken directly from their open-source repository.
Eureka
Eureka’s training-launch success rate was evaluated on their only legged robot environment, ANYmal, across training-launch attempts, successfully launching training times. We also tested training-launch success across all embodiments provided in their open-source repository, which include the Gymnasium environments Shadow Hand, Franka Emika Panda, Ant, Humanoid, and Cartpole, as well as other robot environments such as ANYmal, Allegro, Ball Balance, and Quadcopter.
-B AURA LLM Agent Prompt Templates
The prompts below define the instruction templates used by AURA’s LLM agents throughout the curriculum generation and training process. These templates structure how input data, such as user task prompts, stage descriptions, and past training artifacts, are transformed into schema-compliant outputs like workflows, configuration files, and reward functions. Placeholders in the form of <INSERT_...> denote dynamic inputs that are automatically populated by the AURA framework at runtime with the appropriate content, ensuring that each prompt remains generalizable while retaining full contextual relevance for the target task.
-B1 Curriculum LLM
-B2 RAG
These prompts are used within the RAG Block to query, retrieve, and select suitable past workflows and training files to be used as context for the current task.
VDB Query LLM
Selector LLM
-B3 Stage-Level LLM
-B4 Feedback LLM
The above feedback prompt, together with the next stage’s description and all the stage’s metrics log file, is prepended to the Per-Stage LLM prompt for use by the Feedback LLM.
-C AURA Outputs
Configuration YAML
Reward YAML
Randomization YAML
Reward YAML
Randomization YAML